OpenStudy (anonymous):

Fan&Medal Which expression is the simplified form of the square root of k to the seventeenth power.? k8the square root of k. kthe square root of k to the fourth power. k4the square root of k. k5the square root of k to the seventh power.

OpenStudy (gorv):

first one

OpenStudy (gorv):

k^8*k^1/2 \[\sqrt{k^{17}}=\sqrt{k^{16}*k}=k^{8}*\sqrt{k}\]